OpenStudy (anonymous):

2x^2-7=7x use quadratic formula to solve .

OpenStudy (anonymous):

a= 2, b=-7, c=-7 so you get x=1/4(7-sqrt(105)) and also x=1/4(7+sqrt(105))

OpenStudy (anonymous):

after plugging a, b and c into the quadratic eq.

OpenStudy (anonymous):

the quadratic equation is \[\frac{-b \pm \sqrt{b^2-4*a*c}}{2*a}\] in your equation \[2x^2-7=7x\] you need to rearrange it so that the equation is equal to 0, *hint* move 7x to the left side*

OpenStudy (anonymous):

\[2x^2-7x-7=0\], now \[a*x^2 + b*x +c\] is a generic quadratic polynomial, your polynomial is of similar form, and you actually know what a,b,c are equal to. Now you simply need to plug the values in the the quadratic equation